OpenStudy (anonymous):

find the exact value of x in the following 2^x=5

OpenStudy (lgbasallote):

change to log form first \[\huge a^b = c \implies \log_a c = b\] therefore... \[\huge 2^ x = 5 \implies \log_2 5 = x\] make sense?

OpenStudy (anonymous):

its going to have to make sense, just show mw where i can go from here with this to get In5/In2

OpenStudy (lgbasallote):

\[\huge \log _2 5 \implies \log 5 \div \log 2 \implies \frac{\log 5}{\log 2}\]\[\huge \implies \frac{\ln 5}{\ln 2}\]

OpenStudy (anonymous):

thanks for that, i'm learning calculus and you've been the best help so far, none of the text books i have on calculus contain anything on surds indices or logarithms should i be looking elsewhere for information on these equation types

OpenStudy (lgbasallote):

this is calculus? or precalculus?

OpenStudy (anonymous):

pre calculus i guess

OpenStudy (lgbasallote):

hmm well regarding your question i suggest you research on it,..especially logarithms.because you use logarithms a lot in calculus

OpenStudy (anonymous):

i'll do that tomorrow but until then i'll keep posting my queries and hope you come to the rescue again

OpenStudy (lgbasallote):

sure